# Energy Matching: Unifying Flow Matching and Energy-Based Models for Generative Modeling

### v1

@@ 3.3 Local intrinsic dimension estimation
Real-world datasets, despite displaying a high number of variables, they can often be represented by a lower-dimensional manifold—a concept referred to as the manifold hypothesis (Fefferman et al., 2016). The dimension of this manifold is called the intrinsic dimension. Estimating the LID (Vapnik, 1995) at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities. We defer the precise definition and additional details to Appendix B.
@@
Unlike diffusion models, EBMs explicitly parametrize the relative data likelihood. This explicit parametrization enables efficient analysis of the curvature of the underlying data manifold – in this example, estimating the LID. To this end, we compute the Hessian matrix $\nabla_{x}^{2}V(x_{\mathrm{data}})$ at a given data point and perform its spectral decomposition. We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=10$ for MNIST(Deng, 2012) and $\tau=2$ for CIFAR-10). The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension. As shown in Table 2, the LID estimates we obtain exhibit consistently stronger correlations with PNG compression size [^1] PNG is a lossless compression scheme specialized for images and can be used when no LID ground truth is available Kamkari et al. (2024). (evaluated on 4096 images) using Spearman’s correlation. Figure 2 offers qualitative illustrations. Our EBM-based approach outperforms diffusion-based methods by relying on fewer approximations, since the computation is performed exactly on the data manifold rather than in its proximity.
@@
Energy Matching (Ours) 0.895 0.903
Table 2: We summarize the Spearman’s correlation coefficients of LID estimates with PNG compression rate. Benchmarks results reported in (Kamkari et al., 2024).
Figure 2: Qualitative results for LID estimation using the Hessian spectrum of $V_{\theta}(x)$. Left: Spectrum for a low-LID image. Right: Spectrum for a high-LID image. The eigenvalues quantify curvature along principal directions (eigenvectors). A degenerate spectrum (many near-zero eigenvalues) indicates locally "flat" regions, revealing the dimension of the local data manifold (LID). Intuitively, higher image complexity often corresponds to a higher LID.
@@ 4 Conclusion and limitations
  • Learns a static scalar potential energy whose gradient drives rapid high-fidelity sampling–surpassing state-of-the-art energy based models–while also forming a Boltzmann-like density well suitable for controlled generation. All without auxiliary generators.
  • Offers efficient sampling from target data distributions on par with the state-of-the-art, while learning the score at the data manifold with minimal trainable parameters overhead.
@@
In contrast to the widespread belief, we demonstrated that static curl-free methods for generative flows work and offer an exciting direction for future research. Our Energy Matching approach has the potential to yield novel insights into conditional generation and inverse problems for cancer research Weidner et al. (2024); Balcerak et al. (2024), molecules and proteins Wu et al. (2022); Bilodeau et al. (2022), and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
@@ A.2 Ablation on the sampling time
Here, we provide ablation studies on CIFAR-10 unconditional generation. Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$. This procedure produces a stable Boltzmann distribution from which one can sample. In the case of $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$. The reason for this is that once the samples move close to the data manifold there is no Boltzmann-like potential well to keep them from drifting away. We report results for different $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
@@
Appendix B Details on LID estimation
@@
Appendix C Training Details
Below, we detail the training configurations for CIFAR-10, CelebA, and MNIST. The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd.
@@ Appendix C Training Details
The architecture is shown in Figure 5. We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively static) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output. Hyperparameters are: $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 200k iterations using Algorithm 1 and then 25k more with Algorithm 2 on 4xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$
@@
We scale the CIFAR-10 model by 2$\sim$ 2$\times$ 2 ×; see Figure 6. We keep $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 then 25k with Algorithm 2 on 4xA100. The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=2\times 10^{-5}$.
@@
We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\lambda_{\mathrm{CD}}=0.25$, $\Delta t=0.05$, and $M_{\text{Langevin}}=50$. The architecture consists of a two-stage CNN followed by a two-layer MLP, comprising a total of 2M parameters.
@@
Figure 5: Diagram of our UNet+Transformer EBM for CIFAR-10. A UNet (37M params) processes a 3$\times$32$\times$32 image; its output is fed into a Transformer head (PatchEmbed + 8-layer ViT, 12M params) that produces a scalar potential. Here we employ the identical UNet architecture as in [Tong et al., 2023].

### v3

@@ 3.3 Local intrinsic dimension estimation
Real-world datasets, despite displaying a high number of variables, can often be represented by lower-dimensional manifolds—a concept referred to as the manifold hypothesis (Fefferman et al., 2016). The dimension of such a manifold is called the intrinsic dimension. Estimating the LID at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities. We defer the precise definition to Appendix C.
@@
Unlike diffusion models, EBMs explicitly parametrize the relative data likelihood. This explicit parametrization enables efficient analysis of the curvature of the underlying data manifold – in this example, estimating the LID. To this end, we compute the Hessian matrix $\nabla_{x}^{2}V(x_{\mathrm{data}})$ at a given data point and perform its spectral decomposition. We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=10$ for MNIST (Deng, 2012) and $\tau=2$ for CIFAR-10). The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension. As shown in Table 3, the LID estimates we obtain exhibit stronger correlations with PNG compression size [^2] PNG is a lossless compression scheme specialized for images and can provide useful guidance when no LID ground truth is available Kamkari et al. (2024). (evaluated on 4096 images) using Spearman’s correlation. Figure 4 offers qualitative illustrations. Our EBM-based approach compares favorably to diffusion-based methods, as it relies on fewer approximations by performing computations exactly on the data manifold rather than merely in its vicinity.
@@
Energy Matching (Ours) 0.895 0.901
Table 3: Spearman’s correlation coefficients of LID estimates with PNG compression rate. Benchmarks results reported in (Kamkari et al., 2024).
Figure 4: Qualitative results for LID estimation using the Hessian spectrum of $V_{\theta}(x)$. Left: Spectrum for a low-LID image. Right: Spectrum for a high-LID image. The eigenvalues quantify curvature along principal directions (eigenvectors). A degenerate spectrum (many near-zero eigenvalues, marked in red) indicates locally "flat" regions, revealing the LID. Intuitively, higher image complexity often corresponds to a higher LID.
@@ 4 Conclusion and limitations
  • Learns a time-independent scalar potential energy whose gradient drives rapid high-fidelity sampling–surpassing state-of-the-art energy-based models–while also forming a Boltzmann-like density well suitable for controlled generation. All without auxiliary generators.
  • Offers efficient sampling from target data distributions on par with the state-of-the-art, while learning the score at the data manifold with manageable trainable parameters overhead.
@@
In contrast to the widespread belief, we demonstrated that time-independent curl-free methods for generative flows work and offer an exciting direction for future research. Our Energy Matching approach has the potential to yield novel insights into controlled generation and inverse problems for cancer research Weidner et al. (2024); Balcerak et al. (2024), molecules and proteins Wu et al. (2022); Bilodeau et al. (2022), computational fluid dynamics Gao et al. (2024); Shysheya et al. (2024); Molinaro et al. (2024) and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
@@ A.2 Ablation on the sampling time
Here, we provide ablation studies on CIFAR-10 unconditional generation. Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, producing a stable Boltzmann distribution from which one can sample. Figure 6 illustrates the FID as a function of sampling time $\operatorname{\tau_{s}}$ for models trained under these different regimes. In the case of pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$; this occurs because, once the samples move close to the data manifold, there is no Boltzmann-like potential well to keep them from drifting away. Because the fidelity slope near the data manifold is steep with respect to sampling time, methods lacking explicit time-conditioning can easily overshoot or undershoot, significantly impacting fidelity. This behavior might explain why some models degrade in performance when made time-independent, as recently reported by Sun et al. [2025].
In Figure 6 we also report results for different values of the temperature-switching parameter $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
@@
Appendix B Posterior sampling
This section provides additional algorithmic details and experimental settings for inverse problem sampling tasks presented in the main text.
B.1 Algorithm description
The detailed sampling procedure for inverse problems, optionally incorporating an interaction energy term to explicitly control the diversity among generated samples, is provided in Algorithm 3. The Langevin dynamics alternates between evaluating a composite potential $U_{\theta}(x)$, composed of a data fidelity term, a learned energy-based prior $V_{\theta}$, and, optionally, an interaction energy $W(x,x^{\prime})$. This interaction term can, for instance, encourage diversity among parallel chains. Samples are initialized independently from the Gaussian prior and iteratively refined through gradient-based updates. Specific hyperparameters, including the discretization step size $\Delta t$, total sampling time $T$, and the interaction energy strength, are detailed in Appendix D.
1:for $m=1$ to $M$ do
3:end for
5:for $n=0$ to $N-1$ do
7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}-\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
10:    end for
11:end for
12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples
B.2 AAV inverse design protein generation
We consider protein fitness optimization for the adeno-associated virus (AAV) in the medium and hard data regime as proposed by [Kirjner et al., 2024]. The setup for the latent encoding as well as the backbone architecture for the learned prior in the latent space follows [Bogensperger et al., 2025]. Conditional sampling uses classifier guidance via learned predictor networks $g_{\phi}$ or $\tilde{g}_{\phi}$ to guide samples toward high-fitness regions. The CNN-based networks predicting fitness for AAV sequences are directly sourced from [Kirjner et al., 2024] and are trained only on the limited training data for the medium and hard setup, respectively.
For training we follow Algorithm 1 and Algorithm 2. We then sample 128 sequences using Algorithm 3 (same batch size for all baselines), the hyperparameters are detailed in Appendix D. Generated sequences are evaluated for fitness using the learned oracle from Kirjner et al. [2024]. Further, they are assessed for diversity within the generated sequences and novelty relative to the training set [Jain et al., 2022]. As shown in Table 4, our method achieves state-of-the-art fitness while boosting diversity, and incorporating interaction energy in Algorithm 3 further increases diversity with only a minimal impact on fitness.
Table 4: AAV optimization results. For VLGPO (flow-based) and Energy Matching, medium difficulty uses $g_{\phi}$, hard difficulty uses $\tilde{g}_{\phi}$. Metrics (Fitness↑, Diversity↑, Novelty↑). Reported uncertainty of Fitness is expressed as standard deviation.
AAV medium AAV hard
Method Fitness↑ Diversity↑ Novelty↑ Fitness↑ Diversity↑ Novelty↑
Learning Unnormalized Data Likelihood
Energy-based Models
Energy Matching (Ours) 0.59 (0.0) 5.86 5.0 0.61 (0.0) 4.77 6.7
Energy Matching (+repulsion) (Ours) 0.58 (0.0) 6.22 5.0 0.60 (0.0) 5.22 6.6
Learning Transport/Score Along Noised Trajectories
Flow-based Models
VLGPO [Bogensperger et al., 2025] 0.58 (0.0) 5.58 5.0 0.61 (0.0) 4.29 6.2
Diffusion Models
gg-dWJS [Ikram et al., 2024] 0.48 (0.0) 9.48 4.2 0.33 (0.0) 14.3 5.3
Other Methods
LatProtRL [Lee et al., 2024] 0.57 (0.0) 3.00 5.0 0.57 (0.0) 3.00 5.0
GGS [Kirjner et al., 2024] 0.51 (0.0) 4.00 5.4 0.60 (0.0) 4.50 7.0
AdaLead [Sinai et al., 2020] 0.46 (0.0) 8.50 2.8 0.40 (0.0) 8.53 3.4
CbAS [Brookes et al., 2019] 0.43 (0.0) 12.70 7.2 0.36 (0.0) 14.4 8.6
GWG [Grathwohl et al., 2021] 0.43 (0.1) 6.60 7.7 0.33 (0.0) 12.0 12.2
GFN-AL [Jain et al., 2022] 0.20 (0.1) 9.60 19.4 0.10 (0.1) 11.6 19.6
Appendix C Details on LID estimation
@@
Appendix D Training details
Below, we detail the training configurations for CIFAR-10, ImageNet 32x32, CelebA, MNIST, and AAV. The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd [Ansel et al., 2024]. We optimize all models with the Adam optimizer Kingma and Ba [2014], and for CIFAR-10, ImageNet 32x32, and CelebA we maintain an exponential moving average (EMA) of the weights.
While we specifically adopt (i) a one-sided trimmed mean of negative sample energies and (ii) clamping of the contrastive loss for stability, any commonly used EBM technique (e.g., persistent contrastive divergence [Tieleman, 2008], replay buffers, multi-scale negative sampling) could be readily employed.
In our approach, we introduce two hyperparameters, $\alpha$ and $\beta$, to control these stabilizing techniques:
[…]
$\beta$ $=\text{clamp threshold for $\mathcal{L}_{\mathrm{CD}}$ (i.e., we clamp $\mathcal{L}_{\mathrm{CD}}$ to be $\geq-\beta$).}$.
@@ Appendix D Training details
The architecture is shown in Figure 7. We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively time-independent) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output. Hyperparameters are: $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 200k iterations using Algorithm 1 with EMA $0.9999$ and then 10k more with Algorithm 2 and EMA $0.999$ on 4xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=8\times 10^{-3},\alpha=0.1,\beta=-0.05$.
ImageNet 32x32.
The architecture is shown in Figure 7 (same as for CIFAR-10). Hyperparameters are: $\tau_{s}=1.5$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 600k iterations using Algorithm 1 with EMA $0.9999$ and then 7k more with Algorithm 2 and EMA $0.999$ on 7xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$ $\alpha=0.2,\beta=-0.01$.
@@
We scale the CIFAR-10 model by 2$\sim$ 2$\times$ 2 ×; see Figure 8. We keep $\tau_{s}=2.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 with EMA $0.9999$ then 7k with Algorithm 2 and EMA $0.999$ on 4xA100. The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
@@
We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\Delta t=0.05$, $M_{\text{Langevin}}=50$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=0.25$. The architecture consists of a two-stage CNN followed by a two-layer MLP, comprising a total of 2M parameters.
AAV.
We adopt the one-dimensional CNN architecture as used in [Bogensperger et al., 2025], summing the final-layer activations to obtain the potential. We train for 10k iterations using Algorithm 1 and for 1k iterations using Algorithm 2 on a single GeForce RTX 4090. The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$. For Algorithm 3 we use $\tau_{s}=1.7$ for AAV medium and $\tau_{s}=1.3$ for AAV hard, $\tau^{*}=0.9$, $\zeta=0.01$ for AAV medium and $\zeta=0.009$ for AAV hard and $\Delta t=0.01$. We set the target fitness to $y=1$ to aim for the maximum fitness in the generated sequences.
@@
Figure 7: Diagram of our UNet+Transformer EBM for CIFAR-10 and ImageNet 32x32. A UNet (37M params) processes a 3$\times$32$\times$32 image; its output is fed into a Transformer head (PatchEmbed + 8-layer ViT, 12M params) that produces a scalar potential. Here we employ the identical UNet architecture as in [Tong et al., 2023], but with the time parameter fixed at $t=0$ to render the model time-independent.